OpenStudy (anonymous):

Stuck! Please help! "What is the domain and range of the rainbow (parabola f(x) -x^2 + 36) Explain what the domain and range represent. Do all of the values make sense in this situation? Why or why not?" I'm not completely sure what they're asking for? I'm sure the answer is really simple but these questions are very vague to me...

OpenStudy (usukidoll):

domain lies in the x axis and range lies in the y axis

OpenStudy (usukidoll):

is your parabola -x^2+36 or -(x^2+36) -x^2-36?!!!

OpenStudy (anonymous):

-x^2 + 36

OpenStudy (anonymous):

I understand what domain and range is, I'm just not completely sure about what they're asking for when they say what do the domain and range represent in the rainbow and if all the values make sense in the situation

OpenStudy (anonymous):

How would I answer "Do all of the values make sense in this situation?"

OpenStudy (idku):

f(x) = -x^2 + 36 that is what you meanc ?

OpenStudy (idku):

We say "the domain is x" what this means is that the domain is any number that you can plug in for x (without getting undefined output) is the domain.

OpenStudy (idku):

We say "the range is y" this means that range is any value of y that you can possibly get in a function.

OpenStudy (idku):

For example, your case: you have \(\large f(x)=-x^2-36\) is there a number that you can't plug for x because are going to get an undefined output, or there is no such a number?

OpenStudy (idku):

oh, \(f(x)=-x^2+26\) but for domain that doesn't matter

OpenStudy (idku):

quoting: "what do the domain and range represent in the rainbow and if all the values make sense in the situation " that I do not represent either...

OpenStudy (idku):

what a silly question they asked-:(

OpenStudy (anonymous):

I know. The only part I'm really stuck on is if all the values make sense in the situation

OpenStudy (anonymous):

The domain of f(x) is (-∞, ∞), but all values of x don't make sense for the application. For example, if you were to consider the x-axis as the ground (or whatever surface the rainbow starts and ends on), then the domain for this application is [-6, 6]. Similarly the range of f(x) is (-∞, 36), but only [0, 36] make sense.

OpenStudy (anonymous):

what is the range and domain if you dont mind me asking

OpenStudy (anonymous):

@ruby71

OpenStudy (anonymous):

@idku

OpenStudy (usukidoll):

@kimberlylopez the domain is usually on the x-axis (left to right) and the range is usually on the y-axis (down to up). I agree with @peachpi that the domain is indeed from -6 to 6 ... [-6,6]... the range is negative infinity to 36 but only 0 to 36 makes sense..